OpenStudy (anonymous):

Is my answer true? \[\frac3{5i}=\frac{3i}{5}\]

OpenStudy (anonymous):

or does i*i=-1? and why would it =-1?

OpenStudy (anonymous):

@zepdrix

OpenStudy (slaaibak):

No, your answer is false. i*i = -1

OpenStudy (slaaibak):

\[i = \sqrt{-1} \rightarrow i^2 = - 1\]

OpenStudy (anonymous):

ahhhh! yes! that makes sense!!!!

OpenStudy (anonymous):

thanks!